OpenStudy (anonymous):

find the surface area of the part of the cone z=sqrt(x^2+y^2) that lies between the plane y=x and cylinder y = x^2

OpenStudy (anonymous):

Have you already tried parametrizing your surface with polar coordinates (cylindrical specifically)?

OpenStudy (anonymous):

I don't really understand this stuff. Teacher makes mistakes in his examples of it which just confuses me.I honestly don't know where to start.

OpenStudy (anonymous):

.....I did try one method kept getting the double integral of sqrt(2) and had no idea what the region I was integrating on was.

OpenStudy (anonymous):

Sorry, but I don't think that the graph of y=x^2 rotated around the y-axis is a cylinder...

OpenStudy (anonymous):

Sounds like you're attached to circular cylinders, Ti. \(y=x^2\) is a perfectly good parabolic cylinder.

OpenStudy (anonymous):

its a 3d graph, not a rotation

OpenStudy (anonymous):

I'd recommend using the first formula on this page: http://tutorial.math.lamar.edu/Classes/CalcIII/SurfaceArea.aspx and then trying a polar transform.

OpenStudy (anonymous):

Slaw I tried parametrizing with x=x y=x and x=f(x)=sart(x^2+y^2) and used the formula \[\int\limits_{?}^{?}\int\limits_{?}^{?}\sqrt{1+f _{x}^2+f_{y}^2}dA\]

OpenStudy (anonymous):

when I got the double integral of sqrt 2

OpenStudy (anonymous):

whats the region D then?...are they from x^2 to x and 0 to 1?

OpenStudy (anonymous):

I don't trust myself here cause I don't fully understand the concept.....

OpenStudy (anonymous):

It's the region in the \(x,y\)-plane bounded by those two curves.

OpenStudy (anonymous):

so its \[\int\limits\limits_{0}^{1}\int\limits\limits_{x^2}^{x}\sqrt{1+f _{x}^2+f_{y}^2}dA\] in this case?

OpenStudy (anonymous):

More formally, \[D=\{(x,y)|0 \leq x\leq 1, x \leq y \leq x^2\}.\] Which gives rise to precisely the bounds you have, yes.

OpenStudy (anonymous):

thank you.....its hard to tell if I'm doing it right when the professor screws up the boundaries in class

OpenStudy (anonymous):

I know the feeling. Try deriving that formula using a parametrization; it really helped me understand it when I first learned this stuff.